B & N Ltd monthly demand for shirts is given by p=1500-0.4q,where p denotes the price (in dollars) of each shirt and q denotes the quantity demanded.The company's manufacturing department estimated the total profit (in dollars) of producing q units of shirts to be P (q)=-0.4q2+1300q+1000,Calculate the price at which Buck&MaraghLtd.profit maximized.


1
Expert's answer
2021-11-16T11:12:28-0500

profit maximized if

"P' (q)=-0.8q+1300=0"

"q=1300\/0.8=1625"


"p(1625)=1500-0.4\cdot 1625=\$850"